OpenStudy (anonymous):

What is the period of the sinusoid given by -4sin(2π/(3)x)?

OpenStudy (anonymous):

\[2\pi\] Maybe as it is sin afterall

OpenStudy (anonymous):

Looking for the period of the function below ? \[-4\sin(\frac{2\pi}{3}x) \]

OpenStudy (anonymous):

The period of sin(a) is 2*pi in terms of a in our case a=(2*pi/3)x so the period of sin() in terms of x will be 3